OpenStudy (anonymous):

can i have like some help here O.o -.- okay whats the pattern 1,2,4,7,11,16, what else whats the pattern please help -.-

OpenStudy (anonymous):

2= previous(1)+1 ...4=previous(2) + 2 ...7=previous(4)+3 11=previous(7) +4